OpenStudy (anonymous):

The sum of two consecutive integers is 59. Write an equation that models this situation and find the values of the two integers.

OpenStudy (imstuck):

If we call the first integer x, then what do we have to do to x to get to the next number right after it?

OpenStudy (anonymous):

so you have two integers. Let's call them a and b. \[a+b=59\] but you also know that b is the next number between 0 and 59 after a \[b=a+1\] now use the second equasion by substituting the value of b in the first equasion a+a+1 = 59 simplify and calculate a = ? Once you know a, you know b because a+1 = b = ?

OpenStudy (anonymous):

you see the one of the problems i'm not sure

OpenStudy (imstuck):

I just used x for the first integer, and then x + 1 for the next. The sum of these two is represented as: x + (x + 1) = 59. x + x + 1 = 59... 2x + 1 = 59. Can you do that math?

OpenStudy (anonymous):

consecutive integers means two integers where the second is one higher than the first x + (x+1) or a+b where b=a+1

OpenStudy (anonymous):

yes I can and i've been but to it but its like non of these answers add up to it these ones I have to chose from

OpenStudy (anonymous):

What do you have as a calculated result, and what are the options?

OpenStudy (anonymous):

I have 4 to chose from 1: n+n+1=59;n20;+1=31 2: n+n2=59;n19;2n=38 3: n+n+1=59;n=29;n+1=28 4: n+n+1=59;n=29;n+1=30

OpenStudy (anonymous):

n+n+1 is the correct expression, so 2 is wrong already if you work out n+n+1=59 -> 2n+1 = 59 -> 2n = 59- 1 -> 2n=58 -> n= 29 so n+1= 30 the last option is the correct one.

OpenStudy (anonymous):

so the answer would be 4

OpenStudy (anonymous):

like the 4th one

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you so much!!